XYZ Inc. sells two products, A and B. The selling price, variable cost and contribution margin per unit of product are shown below: Product A B Selling Price $24 $28 Variable Cost $10 $24 Contribution Margin $14 $4 Fixed costs amount to $640,000. The company sells twice as many units of A as it does B. How many units of each product does the company have to sell to breakeven?

Multiple Choice

  • A: 10,000 units; B: 5,000 units.

  • B: 9,000 units; B: 4,500 units.

  • C: 12,000 units; B: 6,000 units.

  • D: 40,000 units; B: 20,000 units.
Answer #1

Given, the company sells twice as many units of A as it does B

If Company sells X units of B, it sells 2X units of A

Then the breakeven units is as follows

[No.of Units of A * Contribution Margin of A] + [ No.of Units of B * Contribution Margin of B] = Fixed Costs

[2X *$14] + [X*$4] = $640,000

28X+4X = $640,000

32X = $640,000

X = $20,000

XYZ inc sells product B of X units = 20,000 Units and product A of 2X units = 2 * 20,000 = 40,000 Units

Hence option D 40,000 units and 20,000 units is correct